# Lepton Number Violation Higgs Decay at Muon Collider

### v1

@@ VI Conclusion
For the heavy Higgs decay signature $H\to{\mu^{\pm}}{\mu^{\pm}}JJ$, a larger heavy Higgs mass usually requires a larger Higgs mixing parameter to reach the 5 $\sigma$ discovery limit. For example, at the 3 TeV muon collider with an integrated luminosity of $L=1~{}\text{ab}^{-1}$, we may discover $m_{H}\lesssim 1$ TeV with $m_{N}=m_{H}/3$ for $\sin\alpha=0.15$. Meanwhile the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature can be detected with the Higgs mixing parameter as small as $\sin\alpha\simeq 0.03$ when $m_{N}=100$ GeV. With $1~{}\text{ab}^{-1}$ data, a large parameter space of the signature $H\to{\mu^{\pm}}{\mu^{\pm}}JJ$ can be detected for $230~{}{\rm{GeV}}<m_{H}<880~{}\rm{GeV}$ and $m_{N}<340~{}\rm{GeV}$ with $\sin\alpha=0.1$. At the 10 TeV muon collider with an integrated luminosity of $L=10~{}\text{ab}^{-1}$, we may discover $m_{H}\lesssim 2.85$ TeV with $m_{N}=m_{H}/3$ for $\sin\alpha=0.15$, meanwhile the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature can be detected with the Higgs mixing parameter down to $\sin\alpha\simeq 0.005$ when $m_{N}=100$ GeV. With $10~{}\text{ab}^{-1}$ data, a large parameter space of the signature $H\to{\mu^{\pm}}{\mu^{\pm}}JJ$ can be detected for $220~{}{\rm{GeV}}<m_{H}<2800~{}\rm{GeV}$ and $m_{N}<1100~{}\rm{GeV}$ with $\sin\alpha=0.1$

### v2

@@ VI Conclusion
For the heavy Higgs decay signature $H\to{\mu^{\pm}}{\mu^{\pm}}JJ$, a larger heavy Higgs mass usually requires a larger Higgs mixing parameter to reach the 5 $\sigma$ discovery limit. For example, at the 3 TeV muon collider with an integrated luminosity of $L=1~\text{ab}^{-1}$, we may discover $m_{H}\lesssim 1$ TeV with $m_{N}=m_{H}/3$ for $\sin\alpha=0.15$. Meanwhile the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature can be detected with the Higgs mixing parameter as small as $\sin\alpha\simeq 0.03$ when $m_{N}=100$ GeV. With $1~\text{ab}^{-1}$ data, a large parameter space of the signature $H\to{\mu^{\pm}}{\mu^{\pm}}JJ$ can be detected for $230~{\rm{GeV}}<m_{H}<880~\rm{GeV}$ and $m_{N}<340~\rm{GeV}$ with $\sin\alpha=0.1$. At the 10 TeV muon collider with an integrated luminosity of $L=10~\text{ab}^{-1}$, we may discover $m_{H}\lesssim 2.85$ TeV with $m_{N}=m_{H}/3$ for $\sin\alpha=0.15$, meanwhile the $H\to\mu^{\pm}\mu^{\pm}JJ$ signature can be detected with the Higgs mixing parameter down to $\sin\alpha\simeq 0.005$ when $m_{N}=100$ GeV. With $10~\text{ab}^{-1}$ data, a large parameter space of the signature $H\to{\mu^{\pm}}{\mu^{\pm}}JJ$ can be detected for $220~{\rm{GeV}}<m_{H}<2800~\rm{GeV}$ and $m_{N}<1100~\rm{GeV}$ with $\sin\alpha=0.1$.